# Refinable modeling for unbinned SMEFT analyses

### v1

@@ 8.8 Results
Figure 11 shows the Asimov expected exclusion contours at 68% CL (dashed) and at 95% CL (solid). The SMEFT effects are simulated up to the quadratic order in the POIs. The other three Wilson coefficients are profiled for the blue contours, while all nuisance parameters are kept at zero. For the black contours, the nuisance parameters are profiled as well. The impacts of the systematic uncertainties are sizable, especially for $\textrm{Re}[C_{tG}]$ which has a sizable effect on the total yield, which is also affected by the luminosity uncertainty, the renormalization and factorization scale uncertainties, and the normalization uncertainties. For non-zero $\textrm{Im}[C_{tG}]$ and $\textrm{Re}[C_{tG}]\approx 0$, the effects of the three remaining four-fermion operators are similar but more significant than the systematic uncertainties, which explains the relatively small extra degradation of the exclusion contours when including systematic uncertainties. Again, we emphasize that a comprehensive assessment of the sensitivity in the $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ final state must include a complete set of uncertainties that, in some cases, are not publicly available. Nevertheless, the study demonstrates how the systematic effects can be cast in machine-learned surrogates and employed in the limit setting.
@@ 9 Conclusion
This paper presents a comprehensive framework for modeling systematic effects in unbinned analyses, particularly for obtaining SMEFT constraints. We factorize the various effects by organizing unobserved (latent) variables and the corresponding systematic effects into parton, particle, and detector levels and make them amenable to ML. We thus leverage the extensive knowledge gained from binned LHC data analyses while maintaining a flexible approach to unbinned modeling. We addressed the need for robust parametric estimates by developing generic methods for training ML surrogates in simulation. Our methodology facilitates the progressive refinement of unbinned models, accommodating new systematic effects or background contributions without invalidating previously trained surrogates.
A significant technical contribution is the parametric regression tree, an extension of tree-boosting algorithms capable of learning accurate parameterizations of systematic effects. This method provides a robust alternative to neural networks, ensuring systematic dependencies can be learned effectively and integrated into unbinned hypothesis tests.
Our work thus bridges a critical gap in the methodological toolbox for SMEFT analyses. We demonstrate the practical application through a semi-realistic case study of top quark pair production in the dilepton channel, which underscores the effectiveness of our approach in learning and incorporating systematic effects. Overall, the techniques and tools developed in this paper pave the way for more refined and adaptable unbinned hypothesis tests, enhancing the accuracy and reliability of SMEFT analyses. We anticipate these advancements will be instrumental in exploiting future collider event simulations for unbinned data analyses.
@@ Acknowledgements
I would like to thank Suman Chatterjee, Claudius Krause, Dennis Schwarz, Nick Smith, and Nicholas Wardle for many useful discussions.
The computational results were obtained using the Vienna Bio Center and the Austrian Academy of Sciences CLIP cluster at https://www.clip.science/.
@@ Appendix B Alternative loss functions
where $L_{1}$ and $L_{2}$, typically, are simple functions of $\hat{f}$. Because it is a sum of expectations over the joint space, this general form allows using the joint-likelihood-ratio in the same way as done for Eq. 44. Moreover, because $\hat{f}(\boldsymbol{x})$ does not depend on $\boldsymbol{z}$, it is minimized by a function of the ratio of two $\boldsymbol{z}$-integrals (Eq. 12) and, therefore, is in one-to-one correspondence with the regression target [^6] I thank Giuliano Panico for pointing this out to me.. If we view the two terms $L_{1}$ and $L_{2}$ as (standard) functions of $\hat{f}$ and denote the (standard) derivative by $L^{\prime}$, it is straightforward to show that the conditions

### v3

@@ 7.8 Results
Figure 11 shows the Asimov expected exclusion contours at 68% CL (dashed) and 95% CL (solid). The SMEFT effects are simulated up to quadratic order in the POIs. For the blue contours, the other three Wilson coefficients are profiled, while nuisance parameters are set to zero. For the black contours, nuisance parameters are also profiled. Systematic uncertainties significantly impact $\textrm{Re}[C_{tG}]$, which strongly affects the total yield and is sensitive to integrated luminosity, renormalization and factorization scales, and normalization uncertainties. For non-zero $\textrm{Im}[C_{tG}]$ and $\textrm{Re}[C_{tG}]\approx 0$, effects from the three remaining four-fermion operators outweigh those of systematic uncertainties, explaining why the contours degrade only slightly when including systematics. A comprehensive $\mathrm{t}\overline{\mathrm{t}}(2\ell)$ sensitivity analysis would require analyzing all uncertainties, some of which are not publicly available. This study, however, shows how systematic effects can be captured in machine-learned surrogates and applied in limit setting.
@@ 8 Conclusion
This paper presents a comprehensive, scalable framework for modeling the effects of systematic uncertainties in unbinned analyses of collider data. By factorizing systematic effects across parton, particle, and detector levels, we make them accessible for machine learning. With a highly granular factorization of the various dependencies, we leverage the extensive knowledge gained from binned LHC data analyses and fully capitalize on high-quality Monte Carlo simulation. A flexible approach facilitates the progressive refinement of unbinned models, including but not restricted to applications in SMEFT. It accommodates new systematic effects or background contributions without invalidating previously trained surrogates.
A significant technical innovation introduced is the Boosted Parametric Tree (BPT), an extension of tree-boosting algorithms designed to learn accurate parametrizations of systematic dependencies. BPTs offer a robust and efficient alternative to neural networks for modeling systematic effects, providing reliable surrogate models for complex, high-dimensional parameter spaces in unbinned hypothesis testing.
Our work thus bridges a critical gap in the methodological toolbox for SMEFT analyses, searches for other non-resonant effects beyond the standard model, and similar inference problems. We demonstrate the practical application through a semi-realistic case study of top quark pair production in the dilepton channel, which underscores the effectiveness of our approach in learning and incorporating systematic effects. Overall, the new techniques pave the way for more refined and adaptable unbinned hypothesis tests, enhancing the accuracy and reliability of SMEFT analyses. We anticipate these advancements will be instrumental in exploiting the data from future collider experiments. Finally, we believe that publicly available refined models would be useful for future SMEFT combinations and for providing legacy LHC results.
@@ Acknowledgements
The computational results were obtained using the Vienna Bio Center and the CLIP computing cluster of the Austrian Academy of Sciences at https://www.clip.science/. I am indebted to Suman Chatterjee, Claudius Krause, Tilman Plehn, Dennis Schwarz, Nick Smith, and Nicholas Wardle for many useful discussions.
@@ Appendix B Alternative loss functions
where $L_{1}$ and $L_{2}$, typically, are simple functions of $\hat{f}$. Because it is a sum of expectations over the joint space, this general form allows using the joint-likelihood-ratio in the same way as done for Eq. 44. Moreover, because $\hat{f}(\boldsymbol{x})$ does not depend on $\boldsymbol{z}$, it is minimized by a function of the ratio of two $\boldsymbol{z}$-integrals (Eq. 12) and, therefore, is in one-to-one correspondence with the regression target [^6] I thank Giuliano Panico for pointing this out.. If we view the two terms $L_{1}$ and $L_{2}$ as (standard) functions of $\hat{f}$ and denote the (standard) derivative by $L^{\prime}$, it is straightforward to show that the conditions
@@
Appendix C Construction of the BPT algorithm
This section provides a step-by-step derivation of the BPT algorithm. A summary of the resulting procedures is described in Sec. 5.4.
C.1 Tree-boosting of parametric regressors
It is instructive to discuss boosting for generic non-parametric estimators based on the cross-entropy loss function $L_{\textrm{CE}}[\hat{f}]$ in Eq. 39. After the replacement in Eq. 47, we have
$L[\hat{T}]$ $=\left\langle\textrm{Soft}^{+}(\hat{T}(\boldsymbol{x}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\mathbf{0}}}+\left\langle\textrm{Soft}^{+}(-\hat{T}(\boldsymbol{x}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}}},$ (C.1)
$T^{\ast}(\boldsymbol{x})=\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})\,},$ (C.2)
but instead of obtaining this result in a single fit, we chose a number $B$ of boosting iterations and corresponding learning rates $10<\eta^{(b)}<$ 1 for $b=1,\ldots,B$. We use an additive expansion of $\hat{T}(\boldsymbol{x})$ in terms of the weak learners $\hat{t}^{(b)}(\boldsymbol{x})$. To this end, we iterate the boosting relations
$t^{(b)}{}^{\ast}(\boldsymbol{x})$ $=\operatorname*{arg\,min}_{\hat{t}^{(b)}}L\left[\hat{t}^{(b)}{}(\boldsymbol{x})+\hat{T}^{(b-1)}(\boldsymbol{x})\right],$ (C.3)
$\hat{T}^{(b)}(\boldsymbol{x})$ $=\hat{T}^{(b-1)}(\boldsymbol{x})+\eta^{(b)}\hat{t}^{(b)}{}^{\ast}(\boldsymbol{x})$ (C.4)
a number of $B$ times, starting with the initial choice $\hat{T}^{(0)}(\boldsymbol{x})=0$. Equation C.3 obtains the weak learner $\hat{t}^{(b)}(\boldsymbol{x})$ when the result of the preceding iteration $\hat{T}^{(b-1)}(\boldsymbol{x})$ is known. Equation C.4 updates the additive model with a fraction $\eta^{(b)}$ of this weak learner’s prediction. After $B$ iterations, the boosted prediction $\hat{T}^{(B)}$ can be expressed as
$\hat{T}^{(B)}(\boldsymbol{x})$ $=\sum_{b=1}^{B}\eta^{(b)}t^{(b)}{}^{\ast}(\boldsymbol{x}).$ (C.5)
An important practicality for boosting learners that are fit to synthetic data sets follows from the minimum condition in Eq. C.3. It implies that the minimum at iteration $b$ satisfies
$t^{(b)}{}^{\ast}(\boldsymbol{x})+\hat{T}^{(b-1)}(\boldsymbol{x})\simeq\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})}=\log\frac{\sigma({\boldsymbol{\nu}})}{\sigma({\mathbf{0}})}\frac{\int{\textrm{d}}\boldsymbol{z}\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})}{\int{\textrm{d}}\boldsymbol{z}\,p(\boldsymbol{x},\boldsymbol{z}|{\mathbf{0}})}$ (C.6)
which we rearrange to
[…]
By reading this equation as an $\boldsymbol{x}$-dependent scaling of the joint-space integration measure ${\textrm{d}}\sigma(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})=\sigma({\boldsymbol{\nu}})\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}}){\textrm{d}}\boldsymbol{x}{\textrm{d}}\boldsymbol{z}$ by the reciprocal of the estimate of the preceding boosting iteration, we find that $t^{(b)}{}^{\ast}$ can also be obtained if, instead of using the additive expansion, we replace $\sigma({\boldsymbol{\nu}})\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})\rightarrow\exp\left(-\hat{T}^{(b-1)}(\boldsymbol{x})\right)\,\sigma({\boldsymbol{\nu}})\,p(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})$. Because Eq. C.4 provides the exponent $\hat{T}$ iteratively, we only have to multiply the cross-section by $\exp(-\eta^{(b-1)}t^{(b-1)}{}^{\ast}(\boldsymbol{x}))$ when moving from iteration $b-1$ to iteration $b$. This way, the boosting equations read
$\sigma({\boldsymbol{\nu}})p^{(b)}(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})$ $=\exp\left(-\eta^{(b-1)}t^{(b-1)}{}^{\ast}(\boldsymbol{x})\right)\,\sigma({\boldsymbol{\nu}})p^{(b-1)}(\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}})$ (C.8)
$t^{(b)}{}^{\ast}(\boldsymbol{x})$ $=\operatorname*{arg\,min}_{\hat{t}^{(b)}}L\left[\hat{t}^{(b)}{}(\boldsymbol{x})\right],$ (C.9)
$\hat{T}^{(b)}(\boldsymbol{x})$ $=\hat{T}^{(b-1)}(\boldsymbol{x})+\eta^{(b)}\hat{t}^{(b)}{}^{\ast}(\boldsymbol{x}),$ (C.10)
initialized by $\hat{T}^{(0)}(\boldsymbol{x})=t^{(0)}{}^{\ast}(\boldsymbol{x})=0$. The advantage of this formulation is that Eq. C.9 is a standard loss function minimization without the additive model appearing in the argument as in Eq. C.3. The update of the synthetic data set $\mathcal{D}_{{\boldsymbol{\nu}}}^{(b)}=\{w_{i}^{(b)},\boldsymbol{x}_{{\boldsymbol{\nu}},i},\boldsymbol{z}_{i}\}$, now also defined for each iteration $b$, follows from Eq. C.8 as
$w_{i}^{(b)}=\exp\left(-\eta^{(b-1)}t^{(b-1)}{}^{\ast}(\boldsymbol{x}_{i})\right)w^{(b-1)}_{i}.$ (C.11)
This prescription can be interpreted as a recursive weighting of the differential cross-section of $\mathcal{D}_{{\boldsymbol{\nu}}}$ in the second term in Eq. C.1 towards $\mathcal{D}_{{\mathbf{0}}}$ in the first term. The boosting algorithm removes the learned approximation from the training data as the regressor learns to approximate the DCR more accurately. It is customary to chose $\eta^{(b)}$ independently of $b$, and values between $10^{-3}$ and $3\cdot 10^{-1}$ for this universal learning rate have proven efficient.
The sample $\mathcal{D}_{{\mathbf{0}}}$ stays unchanged in the boosting procedure because we decided to write the $\boldsymbol{x}$-dependent scaling in Eq. C.8 in the numerator. The choice of only reweighting the sample $\mathcal{D}_{\boldsymbol{\nu}}$ is a critical detail. It holds the key to a boosting algorithm that works for the fully parametric regressor, including the ${\boldsymbol{\nu}}$ dependence. We can construct the loss for a parametric tree-based algorithm from the general parametric loss function in Eq. 51, which is a sum of equally structured terms
$L$ $=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}L_{\textrm{CE}}[\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})]=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left(\left\langle\textrm{Soft}^{+}(\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\mathbf{0}}}+\left\langle\textrm{Soft}^{+}(-\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))\right\rangle_{\boldsymbol{x},\boldsymbol{z}|{\boldsymbol{\nu}}}\right).$ (C.12)
The synthetic data set in the first expectation value in each sum term is always $\mathcal{D}_{\mathbf{0}}$, irrespective of the value of ${\boldsymbol{\nu}}$. The synthetic data set in the second expectation is $\mathcal{D}_{{\boldsymbol{\nu}}}$ and is different for each $\mathcal{{\boldsymbol{\nu}}}\in\mathcal{V}$. It is this term whose synthetic data set changes during the boosting algorithm, and because there is one such set for each ${\boldsymbol{\nu}}\in\mathcal{V}$, the reweighting can be done simultaneously for each ${\boldsymbol{\nu}}$ in the sum over $\mathcal{V}$ in Eq. C.12. Repeating the steps starting at Eq. C.1 with a sufficiently expressive ${\boldsymbol{\nu}}$-dependent function $\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})$, it is straightforward to show that aside from the extra ${\boldsymbol{\nu}}$-dependence in the notation nothing else changes. Concretely, we only need to modify Eq. C.5 to notate the ${\boldsymbol{\nu}}$-dependence in the weak learner $\hat{t}(\boldsymbol{x}|{\boldsymbol{\nu}})$. The other steps follow analogously, and Eq. C.11 generalizes to
$\mathcal{D}_{{\boldsymbol{\nu}}}^{(b)}=\left\{\exp(-\eta^{(b-1)}t^{(b-1)\ast}(\boldsymbol{x}_{i}|{\boldsymbol{\nu}}))\,w^{(b-1)}_{i},\boldsymbol{x}_{i},\boldsymbol{z}_{i}\right\}\;\;\textrm{for all}\;\;\{w^{(b-1)},\boldsymbol{x}_{i},\boldsymbol{z}_{i}\}\in\mathcal{D}_{{\boldsymbol{\nu}}}^{(b-1)}\;\;\textrm{for all}\;\;{\boldsymbol{\nu}}\in\mathcal{V},$ (C.13)
which is the same as Eq. C.11 except for that it is carried out simultaneously for each ${\boldsymbol{\nu}}\in\mathcal{V}$. This completes the boosting algorithm for generic weak learners, and we can proceed with constructing the tree-based implementation.
C.2 Learning the phase-space partitioning
$\mathcal{X}=\bigcup_{j\in\mathcal{J}}\Delta\boldsymbol{x}_{j}\quad\textrm{and}\quad\Delta\boldsymbol{x}_{j}\cap\Delta\boldsymbol{x}_{j^{\prime}}=\emptyset\;\;\longleftrightarrow\;\;j\neq j^{\prime}.$ (C.14)
$\mathbbm{1}_{j}(\boldsymbol{x})=\begin{cases}1&\textrm{if}\;\boldsymbol{x}\in\Delta\boldsymbol{x}_{j}\\ 0&\textrm{otherwise}\end{cases}$ (C.15)
$\hat{t}(\boldsymbol{x}|{\boldsymbol{\nu}})=\sum_{j\in\mathcal{J}}\mathbbm{1}_{j}(\boldsymbol{x})\,\hat{t}_{j}({\boldsymbol{\nu}}).$ (C.16)
Because the $\boldsymbol{x}$-dependence is only in the index function, we can insert Eq. C.16 into Eq. C.12 and use Eq. 43 to carry out the event sums over the synthetic data sets. The result is
$L[\mathcal{J},\hat{t}_{j}]=\sum_{j\in\mathcal{J}}L_{j}[\hat{t}_{j}]=\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\sigma_{j,0}\,\textrm{Soft}^{+}(\hat{t}_{j}({\boldsymbol{\nu}}))+\sigma_{j,{\boldsymbol{\nu}}}\,\textrm{Soft}^{+}(-\hat{t}_{j}({\boldsymbol{\nu}}))\right].$ (C.17)
$\sigma_{j,0}=\sum_{(\boldsymbol{x}_{i},w_{i})\,\in\,\mathcal{D}_{\mathbf{0}}\cap\Delta\boldsymbol{x}_{j}}w_{i}\qquad\textrm{and}\qquad\sigma_{j,{\boldsymbol{\nu}}}=\sum_{(\boldsymbol{x}_{i},w_{i})\,\in\,\mathcal{D}_{\boldsymbol{\nu}}\cap\Delta\boldsymbol{x}_{j}}w_{i}$ (C.18)
Before we tackle these problems, it is instructive to develop an intuition for the loss function in Eq. C.17. We assume an infinitely expressive $\hat{t}_{j}({\boldsymbol{\nu}})$ and functionally differentiate Eq. C.17 to arrive at
$0=\frac{\delta L_{j}}{\delta\hat{t}_{j}}=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\frac{\sigma_{j,{\mathbf{0}}}}{1+\exp\left(-\hat{t}_{j}({\boldsymbol{\nu}})\right)}-\frac{\sigma_{j,{\boldsymbol{\nu}}}}{1+\exp\left(\hat{t}_{j}({\boldsymbol{\nu}})\right)}\right].$ (C.19)
This equation is satisfied exactly if
$\hat{t}_{j}({\boldsymbol{\nu}})=\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\quad\textrm{for all}\quad{\boldsymbol{\nu}}\in\mathcal{V},$ (C.20)
which we can always fulfill for sufficiently expressive $\hat{t}_{j}({\boldsymbol{\nu}})$; a perfect representation of the ${\boldsymbol{\nu}}$-dependence in each tree node $j\in\mathcal{J}$ reproduces the logarithm of the DCR for those values of ${\boldsymbol{\nu}}$ whose synthetic data sets we included in the loss function. The predictive function of a tree is finitely expressive. Hence, we will seek an approximation for Eq. C.20 in the next section. But we can meanwhile use the result to shed light on the loss function Eq. C.17. Formally eliminating $\hat{t}_{j}({\boldsymbol{\nu}})$ in favor of its predictions at the points ${\boldsymbol{\nu}}\in\mathcal{V}$, we express the loss solely in terms of the per-bin cross-sections $\sigma_{j,{\mathbf{0}}}$ and $\sigma_{j,{\boldsymbol{\nu}}}$,
$L[\mathcal{J}]$ $=\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\sigma_{j,{\mathbf{0}}}\log\left(1+\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\right)+\sigma_{j,{\boldsymbol{\nu}}}\log\left(1+\frac{\sigma_{j,{\mathbf{0}}}}{\sigma_{j,{\boldsymbol{\nu}}}}\right)\right].$ (C.21)
This equation would provide a loss function for finding the optimal phase space partitioning if we did not need to use finitely expressive $\hat{t}_{j}({\boldsymbol{\nu}})$. If we assume small ${\boldsymbol{\nu}}$ such that a Taylor expansion of $\sigma_{j,{\boldsymbol{\nu}}}$ around ${\boldsymbol{\nu}}={\mathbf{0}}$ is a good approximation, we get
$L[\mathcal{J}]$ $=-\frac{1}{4}\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\nu_{a}\nu_{b}I_{(ab),j}+\ldots$ (C.22)
$I_{(ab),j}=\frac{1}{\sigma_{j,{\mathbf{0}}}}\frac{\partial\sigma_{j,{\boldsymbol{\nu}}}}{\partial\nu_{a}}\frac{\partial\sigma_{j,{\boldsymbol{\nu}}}}{\partial\nu_{a}}\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}$ (C.23)
C.3 Terminal node predictions
$\hat{t}_{j}({\boldsymbol{\nu}})$ $=\nu_{a}\hat{\Delta}_{a,j}+\nu_{a}\nu_{b}\hat{\Delta}_{ab,j}+\nu_{a}\nu_{b}\nu_{c}\hat{\Delta}_{abc,j}+\cdots=\nu_{A}\hat{\Delta}_{A,j}$ (C.24)
$\hat{\Delta}_{j,a}\approx\frac{\partial}{\partial\nu_{a}}t_{j}({\boldsymbol{\nu}})\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}=\frac{\partial}{\partial\nu_{a}}\log\frac{{\textrm{d}}\sigma(\boldsymbol{x}|{\boldsymbol{\nu}})}{{\textrm{d}}\sigma(\boldsymbol{x}|{\mathbf{0}})}\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}=s_{a}+\frac{\partial}{\partial\nu_{a}}\log\sigma({\boldsymbol{\nu}})\Bigg{|}_{{\boldsymbol{\nu}}={\mathbf{0}}}\quad\textrm{for}\quad\boldsymbol{x}\in\Delta\boldsymbol{x}_{j}.$ (C.25)
The last expression relates $\hat{\Delta}_{j,a}$ to the well-known score vector $s_{a}$, a sufficient statistic for small ${\boldsymbol{\nu}}$ and, therefore, an optimal observable. The log-derivative of the inclusive cross-section in the last term does not depend on the phase-space partitioning and, thus, is irrelevant to the optimization. The algorithm will, therefore, aim to reduce the expectation of the variance of the score in the training sample. This is, by definition, the negative value of the Fisher information matrix, consistent with the interpretation in the preceding section. Depending on the concrete problem and the desired accuracy, the higher-order terms in Eq. C.25 can improve the parametrization for larger values of ${\boldsymbol{\nu}}$.
$\hat{t}(\boldsymbol{x}|{\boldsymbol{\nu}})=\sum_{j\in\mathcal{J}}\mathbbm{1}_{j}(\boldsymbol{x})\left(\nu_{A}\hat{\Delta}_{A,j}\right).$ (C.26)
$\nu_{A}\hat{\Delta}_{A,j}$ $=\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\quad\textrm{for all}\quad{\boldsymbol{\nu}}\in\mathcal{V}$ (C.27)
are overdetermined if $|\mathcal{V}|>N_{\Delta}$. We note that Eq. C.27 has the same form as Eq. C.20 except for the finite expressivity on the l.h.s. We are content with an approximate solution of the per-node parametrization because boosting the weak learner will iteratively reduce the shortcomings either way. Any deficiency of a concrete weak learner will be reduced in the subsequent boosting iteration. For $|\mathcal{V}|<N_{\Delta}$, the training data cannot provide a unique estimate, and more data sets must be obtained.
The simplest approach for approximately solving Eq. C.27 is by minimizing the mean-squared error separately for each node $j\in\mathcal{J}$,
$L_{j,\textrm{MSE}}[\hat{\Delta}]$ $=\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left(\nu_{A}\hat{\Delta}_{A,j}-\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\right)^{2}.$ (C.28)
It is solved by
$\hat{\Delta}_{A,j}=\left[\;\sum_{\nu\in\mathcal{V}}{\boldsymbol{\nu}}{\boldsymbol{\nu}}^{\textrm{T}}\right]_{AB}^{-1}\left[\sum_{\;{\boldsymbol{\nu}}\in\mathcal{V}}{\boldsymbol{\nu}}\log\frac{\sigma_{j,{\boldsymbol{\nu}}}}{\sigma_{j,{\mathbf{0}}}}\right]_{B}.$ (C.29)
The matrix
$V_{AB}=\left[\sum_{\nu\in\mathcal{V}}{\boldsymbol{\nu}}{\boldsymbol{\nu}}^{\textrm{T}}\right]_{AB},$ (C.30)
It is instructive to check that the weak learner appropriately responds to training data that is perfectly, not just approximately, consistent with the polynomial ansatz. If we take constants $\delta_{A}$ and consider a model that predicts $\sigma_{j,{\boldsymbol{\nu}}}=\exp(\nu_{A}\delta_{A})\sigma_{j,{\mathbf{0}}}$ in a given region, we can insert into Eq. C.29 and find $\hat{\Delta}_{A}=\delta_{A}$, confirming that the algorithm learns the exact solution if it has the chance.
To complete the construction of the weak learner, we insert the ansatz Eq. C.26 into Eq. C.17 and get
$L[\mathcal{J}]$ $=\sum_{j\in\mathcal{J}}\sum_{{\boldsymbol{\nu}}\in\mathcal{V}}\left[\sigma_{j,{\mathbf{0}}}\,\textrm{Soft}^{+}\left(\nu_{A}\hat{\Delta}_{j,A}\right)+\sigma_{j,{\boldsymbol{\nu}}}\,\textrm{Soft}^{+}\left(-\nu_{A}\hat{\Delta}_{j,A}\right)\right],$ (C.31)
where $\hat{\Delta}_{A,j}$ are obtained from Eq. C.29 and $\sigma_{j,0}$ and $\sigma_{j,{\boldsymbol{\nu}}}$ from Eq. C.18. The data samples $\mathcal{D}_{\mathbf{0}}$, used for the prediction of $\sigma_{j,{\mathbf{0}}}$ in the first term, can either taken to be the same or statistically independent samples. This loss function is amenable to standard tree algorithms, for example, the CART algorithm or the “Tree Alternate Optimization” (TAO) [55, 56, 57, 58] algorithm, both providing tree structures with a hierarchical selection using the features $\boldsymbol{x}$ and that satisfy the requirements in Eq. C.14. These algorithms proceed by recursively splitting the training data along either axis-aligned (for CART) or linear combinations of the input features (for TAO), reducing the loss at each iteration. The maximum iteration depth and the minimum number of events in each node are hyperparameters that regularize the fit. If no more splits can be performed, the terminal selections (nodes) represent a phase space partitioning $\mathcal{J}$ of the form in Eq. C.14 and the quantities $\hat{\Delta}_{j}$ can be computed from Eq. C.29. The tree then estimates the (log-)DCR as in Eq. C.26. As a function of $\boldsymbol{x}$, the prediction of a single tree changes discontinuously if $\boldsymbol{x}$ traverses a boundary between nodes, and the possibly poor approximation close to the boundaries weakens the learner. Utilizing boosting, i.e., using a sequence of trees in Eq. C.3 and Eq. C.4, we recover the smooth behavior in $\boldsymbol{x}$ of an arbitrarily expressive regressor. Because each tree in the boosted result in Eq. C.5 is parametric in ${\boldsymbol{\nu}}$, so is the final parametric regression tree.
C.4 Algorithm summary
We can combine the steps in Secs. C.1C.3 to summarize the BPT algorithm. It is an iterative fit of a tree-based weak learner with the loss in Eq. C.31 to the residuals of the preceding boosting iteration whose predictions are obtained from Eqs. C.8C.11. Concretely, we start with training data $\mathcal{D}_{0}$ at a reference point and several synthetic data sets associated with model parameters ${\boldsymbol{\nu}}\in\mathcal{V}$. We must have enough data so $V_{AB}$ has full rank. We fit a weak learner using the CART algorithm. At each iteration, the CART algorithm recursively divides the feature space by greedily selecting the dimension and cut value combination that minimizes the loss function. Overfitting is mitigated by enforcing a maximum tree depth and a minimum number of events in each terminal node. We construct new synthetic data from the weak learner’s prediction using Eq. C.11 to replace $\mathcal{D}_{\boldsymbol{\nu}}$. This reweighting procedure brings the samples $\mathcal{D}_{\boldsymbol{\nu}}$ closer to $\mathcal{D}_{0}$ by an amount controlled by the learning rate $\eta$. We iterate the whole procedure $B$ times and obtain the final result from Eq. C.5 as
$\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}})=\log\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})=\sum_{b=1}^{B}\eta^{(b)}\sum_{j\in\mathcal{J}^{(b)}}\mathbbm{1}_{j}(\boldsymbol{x})\nu_{A}\hat{\Delta}^{(b)}_{A,j},$ (C.32)
C.5 An analytic toy example
To illustrate the BPT, we consider an arbitrarily chosen one-dimensional two-parameter model
${\textrm{d}}\sigma(\boldsymbol{x}|\nu_{1},\nu_{2})$ $=N\exp\left(0.25\left(\nu_{1}\sin(x)+\nu_{2}\cos(0.5x)\right)^{2}\right){\textrm{d}}x$ (C.33)
with support $x\in[-\pi,\pi]$. The logarithm of the cross-section is a quadratic polynomial in ${\boldsymbol{\nu}}$ for all $\boldsymbol{x}$, suggesting a perfect fit with a two-parameter parametric tree at quadratic accuracy. For the training, we chose five base points $\mathcal{V}=\{(0.5,0),\,(0,0.5),\,(1,0),\,(0,1),\,(0.5,0.5)\}$ that lead to a full-rank matrix $V$ in Eq. C.30. With a nominal data set at $(\nu_{1},\nu_{2})=(0,0)$, we have six synthetic data sets, each with $5\cdot 10^{5}$ events, sufficient to train the algorithm. We fit $B=100$ boosting iterations and require a maximum tree depth of $4$ and a minimum requirement for the number of events in each terminal node, which is 50 events. The learning rate is set to 0.2 for all boosting iterations.
Figure 12: The DCR for various model parameters for the toy study described in the text. Dashed lines show the true DCR from the training data; solid lines show the prediction obtained from the BPT algorithm. The top (bottom) panel compares after $B=5$ ($B=100$) boosting iterations. The bottom panel also includes new model parameter points not present during training. In this panel, the prediction is drawn after the truth, and the dashed histograms are not separately visible because of the good quality of the fit with $B=100$.